Operations Report – April 29th

Crew 193 Operations Report 29Apr2018

SOL: 7

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10. Soldering Irons
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).
I will repair charging plugs once we have a soldering iron.
One piece suit #10’s fans stopped when I bent over while wearing the suit. The loose connection was determined to be at the positive terminal of the battery; upon inspection, there was a cold solder joint on a terminal connector. While removing tape to access battery terminals, I broke the small wires on negative terminal. Suit #10 is out of service until we have a working soldering iron (should be on the way).

Generator (hours run): 10:45 pm to 7:35 am.

Solar— SOC 57 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 2.5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 455 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134315.1

Toilet tank emptied: yes
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.3

ATV Oil Checked?: Yes

ATV Oil Added?: No

ATV Fuel Used: 2.5 Gallons

# Hours the ATVs were Used today: 2

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda oil is within range, but on low end.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.
Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: Sawhorses moved in from RAM.

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Would like some new crimp-on connectors for suit battery terminals. For example: https://www.mcmaster.com/#7060K38. An assortment would probably be nice for future missions.

Greenhab Report – April 29th

Green Hab Report
Date: 29 Apr 2018

Environmental control: Ambient with door open but only for few hours in the afternoon
Shade cloth: on
Fan: not used today

Average temperatures: 34.6 C
Low: 14.7℃
High: 51.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 8.5 Gallons

Time(s) of watering for crops: Twice: In the morning at 900am and afternoon 445pm.

Morning research observations: None

Changes to research plants: None

Aquaponics: None

Narrative: No significant changes. Plants that were yellowing are turning green again after extra watering.

Support/supplies needed: None

Sol Summary – April 29th

Crew 193 Sol Summary Report – 29APR2018

Sol 7

Summary Title: How Windy Is It? It’s Sooooollllll Windy, the Answer Will Blow You Away!

Author: Dr. Shawna Pandya, HSO

Mission Status: Supra-nominally windy.

Sol Activity Summary:
Today the two EVA teams swapped AM/PM spots to change things up, and will be doing so for the rest of their time here. EVA-AM team explored White Rock Canyon and Kissing Camels Ridge, and a good time was had by all, other than a sprained knee sustained by one of the team, who is now recovering. Meanwhile, EVA-PM team took care of the GreenHab and wo/manned the fort. Upon her return, the Commander took Command of the kitchen and whipped up a delicious feast of pasta alfredo with chicken, spaghetti bolognaise and a hearty vegetable soup to whet her picky crew’s carnivorous, non-dairy and vegetarian predilections, respectively. In the late afternoon, EVA Team 2 took to the mean streets of Mars for a sojourn up North, certain that they would have turn back early due to mounting winds and increasing cloud cover. Alas, the winds of Mars made way for our fearless explorers, who had the time (and view) of their lives up by Copernicus way. In the midst of these EVAs, more scientific data was accrued and solar prominences observed by the science team and astronomer. The Sol will culminate with another hearty Martian spread come dinnertime, followed by some ‘Gold’-en opportunities for crew bonding, by way of ASL, cards and games.

Look Ahead Plan:
The week continues with more EVAs to the Moon by the EVA-AM team, and a PM trek over to Cactus Road with exploration of the near-by sights and sounds, with plans to forge on with the usual science, Hab operations and astronomy.

Anomalies in work:

1. Generator is not charging solar panels overnight – ongoing issue, no changes.
2. Suit #10 has been decommissioned due to aberrant battery contacts.

Weather: Partly overcast, very windy, warm.

Crew Physical Status: One crew member sprained a ligament on EVA and is recovering. They have plans to go on a light excursion EVA (light walking on flat terrain or mostly in vehicle), or plans to stay back if not recovered in the morning.

EVA:

Reports filed:
EVA Report
Astronomy Report

Science Report
Operations Report

HSO Report
GreenHab Report
Sol Summary

Reports to be filed:
GreenHab
Photo & Photo of the Day Report

Commander’s Report
EVA Request

Support Requested:
1. Straps to 2-piece suit neck ring
2. New terminal connectors for Suit #10

Operations Report – April 28th

Crew 193 Operations Report 28Apr2018

SOL: 6

Name of person filing report: David Attig

Non-nominal systems: Generator – sim suit chargers

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 11:00 pm to 8:30 am.

Solar— SOC 70 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 500 gallons

Trailer to Static Pump used – yes

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134250.5

Toilet tank emptied: no
Date last emptied: April 26

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Outreach videos were made today, water transferred from storage tank to static tank. Red wagon assembled for greenhab.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: One piece suits nominal, still work to do on two piece suits. Repair of suit chargers still in progress.

Summary of Hab operations: Transferred water from storage tank to static tank.

Summary of GreenHab operations: Watered twice, plants returning to health. Assembled red wagon placed inside.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: Not used today.

Summary of health and safety issues: One crewmember scraped their hand when going down the stairs (actually fell down the last one).

Questions, concerns and requests to Mission Support: We would like an ice cube tray for the freezer.

Commander's Report – April 28th

Mission Sol 6 (Earth Day April 28th, 2018)

Sols fly by on Mars! Before we knew it, we are at the half-way point in our mission. We took the time today to ponder over what we have accomplished so far and plan the rest of our mission. It being a rest day, we decided to put it to good use and spent the morning creating outreach videos.

Followed it up with a lunch of quinoa, chicken curry and daal.

Our CGA captured a beautiful view of the prominence on the southeast side of the Sun! Cannot wait to take a look at the processed image. Our CE and HSO spent early afternoon putting together the new red wagon as requested by Mission Control.

We plan on taking it easy the rest of the evening. We’ve been learning the American Sign Language from our CGA (who’s hearing impaired), will dedicate some time this evening for the same, a round of "Never Ever Have I" game is on the cards after and might wrap up our evening with a movie, Hidden Figures Or Interstellar. Both interesting. Let’s see what the crew is in mood for.

Thank you for being with us this evening Mission Support!

Signing off for this Sol!

Warm Regards,
Commander, (Gold 00), Crew #193
Mars Desert Research Station

(P.S. Abbreviations Legend: CO – Commander / XO – Executive Officer / CE – Crew Engineer / HSO – Health Science Officer / CGA – Crew Geologist and Astronomer)

Science Report – April 28th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No EVAs took place today.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar prominence during his imaging session this morning.

3. Shortwave Texting on EVA’s: No EVAs took place today.

4. GPS Route Measurements: No EVAs took place today.

5. Waterless dish cleaning: Samples were taken from a dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria are growing on the samples from the first test.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No EVAs took place today.

8. Crew Comfort, Health and Safety on EVA: No EVAs took place today.

Sol 6 Summary Report – April 28th

Sol 6

Summary Title: Stopping to Smell the Roses

Author: Eric Shear, CGA

Mission Status: Perfectly nominal… for now.

Sol Activity Summary: This morning we did a series of short outreach videos per the astronomer’s suggestion. For lunch, we feasted on (yet) another delicious Martian lunch of quinoa and curry chicken courtesy of our commander’s mean cooking skills. The afternoon was spent on ongoing science experiments, finishing the MDRS sign, building the new red wagon for the Greenhouse, and usual Hab operations. Our crew astronomer managed to catch a small prominence on the sun’s limb during his observations this morning. The remainder of the day was spent planning out our remaining EVAs. We now have a week’s worth of EVA destinations for both away teams.

Look Ahead Plan: Tomorrow, we will execute the first two EVAs of the new week. In the morning, the American team will head to Zubrin’s Rock and explore White Rock Canyon. Team Canada will follow up at Toothy Ridge in the afternoon.

Anomalies in work:

None

Weather: Cloudy skies in AM and PM. Extremely hot.

Crew Physical Status: Nominal

EVA: None

Reports filed:

Astronomy Report

GreenHab Report

Reports to be filed:

EVA Report

EVA Request

Science Report

Journalist Report

Photo & Photo of the Day Report

Operations Report

Support Requested:

None

Astronomy Report – April 28th

Astronomy Report

Name: Eric Shear

Crew: 193
Date: April 28, 2018

Sky Conditions: Sunny with clouds

Wind Conditions: No wind earlier in morning, windy around noon

Observation 1 Start Time: 0945

Observation 1 End Time: 1015

Observation 2 Start Time: 1145

Observation 2 End Time: 1200

Summary:

Observation 1: The solar flare and filament from yesterday were no longer visible with the camera. No sunspots could be seen with the eye, and the camera could not discriminate any from the granules.

Observation 2: I got a heads-up from Peter Detterline about a prominence on the Sun’s southeast side. It was not visible in the eyepiece at first, but after playing with the tuners it came into view. I attached the Skyris camera, brought the feature into focus, and took a capture.

Photos were taken with the Skyris camera and the best images processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The results are attached.

Objects Viewed: Sun

Problems Encountered: Winds during second observation were such that I could not open the lower shutter without it shaking too much, so I kept it closed. The sun was high enough in the sky by then that I could observe it with just the upper shutter retracted.

GreenHab Report – April 28th

GreenHab Report

Name: Shawna Pandya

Date: 28Apr2018

Environmental control: Ambient with door open

Shade cloth: on

Fan: Used today

Average temperatures: 39.0℃

Low: 15.5℃

High: 52.2℃

Hours of supplemental light: None

Changes to crops: Less yellow with additional watering

Daily water usage for crops: 10 Gallons

Time(s) of watering for crops: Twice: Once at 0900 and again at 1700

Morning research observations: None

Changes to research plants: No.

Aquaponics: None

Narrative: Picked a few tomatoes. Assembled red wagon – the latest ‘Mars Rover’!

Support/supplies needed: None